# Thermodynamic geometry in hadron resonance gas model at real and imaginary baryon chemical potential and a simple sufficient condition for quark deconfinement

### v2

@@ V Scalar curvature
In this section, we investigate the scalar curvature $R$ in the HRG models with and without EVE. Since the calculation of $R$ is difficult at low temperature, we consider the cases only in the region of $T=25\sim 300$ MeV. Note that, since $\gamma$ depends not only on $\mu$ but also on $T$, the relation between two coordinates $(T,\mu)$ and $(\beta,\gamma)$ is not simple and we obtain
@@
at $\mu=0$. Hence, $h_{1}=\phi_{\beta\gamma\gamma}=0$, $h=0$ and $R=0$ when $B(T)=1/v_{\rm B}$. The condition $B(T)=1/v_{\rm B}$ is nothing but the condition when the singularity occurs at $\theta=\pi$. Therefore, $R=0$ when $\mu=0$ and $T=T_{1}=T_{\rm RWL}=0.2103$ GeV. We see that $T_{\rm RWL}$ is special temperature even at $\mu=0$.
@@
We confirm that $g_{11}$ and $g$ are always positive as well as the case at $\mu=0$ when $\mu$ is finite and real. Hence, the line element (LABEL:dl2) is positive definite in this case and the system is thermodynamically stable.
@@
Figure 14 shows the curve on which the $R=0$ is satisfied on the $\mu$-$T$ plane in the HRG model without EVE. When $T$ is large and/or $\mu$ is large, $R$ is positive. At small $\mu$, the pseudo-critical temperature obtained by the $R=0$ criterion is much larger than the LQCD crossover temperature. This result is consistent with the previous studies Castorina et al. (2018, 2019). There is no phase structure in the large $\mu$ region.
Figure 15 is the same as Fig. 14 but the HRG model with EVE is used. In this case, $R<0$ when $T$ is very high and when $T$ is small and $\mu$ is small or very large. In the small $\mu$ region, the lower curve of the $R=0$ criterion is consistent with the LQCD crossover temperature up to $\mu\sim 0.1$ GeV. This result is consistent with the previous study Castorina et al. (2019). The crossover temperature line meets the $|R|$ local minimum curve at $\mu\sim 0.25$ GeV. We have also calculated $R$ in the large $\mu$ region. The critical point (CP) predicted by the LQCD calculation lies just below the right-lower curve of the $R=0$ criterion. This small disagreement indicates that the HRG model with EVE is not applicable beyond the critical point. In the next section, we investigate the limitation of the HRG model with EVE.
@@
Fig. 15: The solid lines represent the curves determined by the condition $R=0$ on the $\mu$-$T$ plane in the HRG model with EVE. The dash-dotted line represents the curve on which $|R|$ has its local minimum for the fixed value of $\mu$ in the region where 0.202 $\mu<0.525 GeV$ and $T<T_{\rm RWL}$. The lower (upper) dashed line represents the curve determined by the condition $n_{\rm b}=1/(2v_{\rm B})$ ($n_{\rm a}=1/(2v_{\rm B}$)). The meaning of the symbols is the same as in Fig. 14.
@@
Figure 19 is the same as Fig. 18 but the HRG model with EVE is used. In this case, $R>0$ in the region of intermediate temperature and $\theta<0.36\pi$, and in the vicinity of the RWL point $(\theta,T)=(\pi,T_{\rm RWL})$. The fine structure of the curve of $R=0$ in the vicinity of the RWL point is shown in Fig. 20. $R$ is negative elsewhere. The curve of $R=0$ near the $T$ axis is connected to the RWL region by the curve of $|R|_{\rm min}$. There are two regions where $g^{\prime}=-g>0$ and $g_{11}^{\prime}=g_{11}>0$. There are also two regions where $g_{11}$ is negative. The curve determined by the condition $|n_{b0}|=B(T)=1/v_{\rm B}$ is the horizontal line which connects the upper zero point of $R$ at $\theta=0$ with the RWL point at $\theta=\pi$.
Fig. 19: The solid line represents the curve determined by the condition $R=0$ on the $\theta$-$T$ plane in the HRG model with EVE. The dash-dotted line represents the curve on which $|R|$ has its minimum for the fixed value of $\theta$ in the region of $0.35<\theta<0.9914$ where the zero point of $R$ does not exists. The square is the RWL point $(\theta,T)=(\pi,T_{\rm RWL})$. The dotted (dash-dot-dotted) line represents the curve determined by the condition $g=0$ ($g_{11}=0$). The dashed line represents the curve determined by the condition $|n_{\rm b0}|=B(T)=1/v_{\rm B}$. Note that $|n_{\rm a0}|=|n_{\rm b0}|$ when $\mu$ is imaginary.
Fig. 20: The solid line represents the curve determined by the condition $R=0$ on the $\theta$-$T$ plane in the HRG model with EVE. The dash-dotted line represents the curve on which $|R|$ has its minimum for the fixed value of $\theta$ in the region of $0.99\leq\theta<0.9914$ where the zero point of $R$ does not exists. The square is the RWL point $(\theta,T)=(\pi,T_{\rm RWL})$.
@@
Figure 22 shows the phase diagram determined by the condition $R=0$ on the $\mu^{2}$-$T$ plane when EVE is present. The curve of $R=0$ near the $T$ axis continues analytically from the real $\mu$ region to the imaginary one and is connected to the RWL region by the curve of $|R|_{\rm min}$. The efficiency of the HRG model with/without EVE can be checked by the calculating the scalar curvature $R$ by the LQCD calculation in the imaginary chemical potential region where there is no sign problem. In the imaginary $\mu$ region, $R$ diverges on the curves on which $g=0$. It seems that there is no counterpart of this singularity in the real $\mu$ region, since this divergence of $R$ due to the vanishing of $g$ is also present in the case without EVE.
@@
VI Limitation temperature of baryon gas model
At $\theta=\pi$, $T_{\rm RWL}$ (=0.2103 GeV) can be regarded as the limitation temperature of the HRG model with EVE, to say more exactly, the limitation temperature of the baryon gas model. This limitation is related with the sign changing of the scalar curvature $R$. However, in QCD (not in the HRG model), it is expected that baryon gas transits to the quark matter at $T=T_{\rm RW}$ which is slightly smaller than $T_{\rm RWL}$. On the other hand, in the region of real $\mu$, the critical point (CP) predicted by the LQCD calculation lies just below the right-lower curve of the $R=0$ criterion. This small disagreement indicates that the HRG model with EVE is not applicable beyond the critical point. As is the case of the RW transition, in the region of real $\mu$ in the QCD phase diagram, it is expected that there is true limitation temperature of the baryon gas, which is slightly lower than the temperature of the $R=0$ criterion in the HRG model with EVE. In this section, we investigate the limitation of baryon gas model within the framework of the HRG model with EVE.
@@ VI Limitation temperature of baryon gas model
Figure 23 shows $T$-dependence of $\chi_{2}^{\rm B}/T^{2}$ at $\mu=0$ in the HRG models with EVE and without EVE. In the HRG model without EVE, $\chi_{2}^{\rm B}/T^{2}$ increases monotonically as $T$ increases. In the HRG model with EVE, it has its maximum at $T_{\rm max}=0.198$ GeV. Due to the additional factor $T^{-3}$, $T_{\rm max}=0.198$ GeV is somewhat lower than $T_{\rm RWL}$. We see that the results in the HRG model with EVE are consistent with those in LQCD up to $T_{\rm max}$. This indicates that $T_{\rm max}$ is the limitation temperature of the baryon gas model.
@@
Figure 24 shows $T$-dependence of $\chi_{2}^{\rm B}/T^{2}$ at $\mu=0.6$ GeV in the HRG models with EVE and without EVE. Again, in the HRG model without EVE, $\chi_{2}^{\rm B}/T^{2}$ increases monotonically as $T$ increases. In the HRG model with EVE, it has two local maxima and one local minimum. The temperature of lower (upper) local maxima of $\chi_{2}^{\rm B}/T^{2}$ is somewhat lower than the temperature of lower (upper) local maximum of $\phi_{\gamma\gamma}/T_{\rm RWL}^{3}$. It is also seen that the lower local maximum of $\chi_{2}^{\rm B}/T^{2}$ is the global maximum. Hereafter, we regard the temperature of the global maximum of $\chi_{2}^{\rm B}/T^{2}$ as the limitation temperature of the baryon gas model.
In two dimensional conformal field theory, using the renormalization group method, it was shown that the effective degrees of freedom decrease as the energy scale of the system decreases Zamolodchikov (1986). It is called the c-theorem. Inversely, it may be natural that the effective degrees of freedom increase as the energy scale such as temperature increases. Without EVE, the effective degrees of freedom (EDOF) of the HRG model increases as $T$ increases. However, EVE reduces EDOF and the dimensionless quantity $\chi_{2}^{\rm B}/T^{2}$ has its maximum at $T_{\max}$ where the two effects are balanced. It is unnatural that EVE overcomes the effect of increase of temperature and makes $\chi_{2}^{\rm B}/T^{2}$ decrease as temperature increases. Hence, we regard the lower $T_{\rm max}$ as the limitation temperature of the HRG model with EVE.
@@
Figure 25 shows the $\mu$-dependence of $T_{\rm max}(\mu)$ ($T_{\rm min}(\mu)$) where $\chi_{2}^{\rm B}/T^{2}$ has its local maximum (minimum) for the fixed value of $\mu$ in the HRG model with EVE. We see that the critical point predicted by LQCD is located almost on the lower curve of $T_{\rm max}(\mu)$. This fact supports the our assumption that the lower $T_{\rm max}(\mu)$ curve represents the limitation temperature of the baryon gas model. Above this temperature, quarks deconfine at least partially.
In Fig. 25 the $\mu$-dependence of $T_{\rm max}^{\prime}(\mu)$ ($T_{\rm min}^{\prime}(\mu)$) where $\phi_{\gamma\gamma}$ has its local maximum (minimum) for the fixed value of $\mu$ and the curve determined by the condition $n_{\rm b}=n_{\rm b0}/2=1/(2v_{\rm B})$ ($n_{\rm a}=n_{\rm a0}/2=1/(2v_{\rm B})$) are also shown. Note that $T_{\rm max}^{\prime}(\mu=0)=T_{\rm RWL}$. The lower (upper) curve of $T_{\rm max}^{\prime}$ almost coincides with the curve of $n_{\rm b}=1/(2v_{\rm B})$ ($n_{\rm a}=1/(2v_{\rm B})$) when $\mu$ is large. We see that the lower $T_{\rm max}(\mu)$ curve is just below the curve of $n_{\rm b}=1/(2v_{\rm B})$. This fact means that $n_{\rm b}(\mu,T_{\rm max}(\mu))<1/(2v_{\rm B})$. When $\mu$ is large, $n_{\rm B}\sim n_{\rm b}$. Hence, we obtain a simple sufficient condition $n_{\rm B}>1/(2v_{\rm B})$ for quark deconfinement at large $\mu$, if we regard the lower $T_{\rm max}(\mu)$ as the limitation temperature of baryons.
Fig. 25: The dashed-dotted (solid) line shows the position of the local maximum of $\chi_{2}^{\rm B}/T^{2}$ ($\phi_{\gamma\gamma}/T_{\rm RWL}^{3}$) at fixed value of $\mu$ in the HRG model with EVE, while the dotted (dash-dot-dotted) line shows the position of its local minimum. The lower (upper) dashed line represents the curve determined the condition $n_{b}=1/(2v_{\rm B}$) ($n_{a}=1/(2v_{\rm B})$). The meaning of the symbols is the same as in Fig. 14.
@@
Fig. 28: The phase diagram in $\mu$-$T$ plane. The square point is the cross point given by Eq. (70). The meaning of the lines and the other symbol is the same as in Figs. 14 and 25.
@@
It is very interesting that this cross point is very close to the critical point predicted by LQCD calculation. The crossover transition line is expected to end on the curve of the limitation of the baryon gas model. The endpoint is nothing but the critical point. Similarly, the curve of the $\chi_{2}^{\rm B}T/\varepsilon_{T}$ maximum in the HRG model should end at the cross point where the curve crosses the curve of the baryon gas limitation. Remember that, at $\mu=0.3$ GeV, the $\chi_{2}^{\rm B}T/\varepsilon_{T}$ maximum curve merges with the LQCD crossover transition line. Hence, this cross point may be regarded as an approximate critical point predicted by the HRG model with EVE, although the model has no explicit dynamics of the phase transition. Fig. 28 shows the phase diagram which is predicted by our HRG model with EVE, if we assume that the curve of the $\chi_{2}^{\rm B}/T^{2}$ maximum gives the limiting temperature of baryons. Above the dashed-dotted line, quarks deconfine at least partially. Note also that, below the dashed line, $n_{\rm b}<1/(2v_{\rm B})$ in the HRG model with EVE.
VII Summary

### v3

@@ V Scalar curvature
In this section, we investigate the scalar curvature $R$ in the HRG models with and without EVE. Since the calculation of $R$ is difficult at low temperature, we consider the cases only in the region of T=0.0250.3T=0.025-0.3 GeV. Note that, since $\gamma$ depends not only on $\mu$ but also on $T$, the relation between two coordinates $(T,\mu)$ and $(\beta,\gamma)$ is not simple and we obtain
@@
at $\mu=0$. Hence, $h_{1}=\phi_{\beta\gamma\gamma}=0$, $h=0$ and $R=0$ when $B(T)=1/v_{\rm B}$. The condition $B(T)=1/v_{\rm B}$ is nothing but the condition when the singularity occurs at $\theta=\pi$. Therefore, $R=0$ when $\mu=0$ and $T=T_{1}=T_{\rm RWL}=0.2103$ GeV. We see that $T_{\rm RWL}$ is a special temperature even at $\mu=0$.
@@
We confirmed that $g_{11}$ and $g$ are always positive as in the case at $\mu=0$ when $\mu$ is finite and real. Hence, the line element (LABEL:dl2) is positive definite in this case and the system is thermodynamically stable.
@@
Figure 14 shows the curve on which $R=0$ is satisfied on the $\mu$-$T$ plane in the HRG model without EVE. When $T$ is large and/or $\mu$ is large, $R$ is positive. At small $\mu$, the pseudo-critical temperature obtained by the $R=0$ criterion is much larger than the LQCD crossover temperature. This result is consistent with the previous studies Castorina et al. (2018, 2019). There is no phase structure in the large $\mu$ region.
Figure 15 is the same as Fig. 14 but the HRG model with EVE is used. In this case, $R<0$ when $T$ is very high and when $T$ is small and $\mu$ is small or very large. In the small $\mu$ region, the lower curve of the $R=0$ criterion is consistent with the LQCD crossover temperature up to $\mu\sim 0.1$ GeV. This result is consistent with the previous study Castorina et al. (2019). The crossover temperature line meets the $|RT^{3}|$ local minimum curve at $\mu\sim 0.25$ GeV. We have also calculated $R$ in the large $\mu$ region. The critical point (CP) predicted by the LQCD calculation lies just below the right-lower curve of the $R=0$ criterion. This small disagreement indicates that the HRG model with EVE is not applicable beyond the critical point. In the next section, we investigate the limitation of the HRG model with EVE.
@@
Fig. 15: The solid lines represent the curves determined by the condition $R=0$ on the $\mu$-$T$ plane in the HRG model with EVE. The dash-dotted line represents the curve on which $|RT^{3}|$ has its local minimum for the fixed value of $\mu$ in the region where 0.202 $\mu<0.525 GeV$ and $T<T_{\rm RWL}$. The lower (upper) dashed line represents the curve determined by the condition $n_{\rm b}=1/(2v_{\rm B})$ ($n_{\rm a}=1/(2v_{\rm B}$)). The meaning of the symbols is the same as in Fig. 14.
@@
Figure 19 is the same as Fig. 18 but the HRG model with EVE is used. In this case, $R>0$ in the region of intermediate temperature and $\theta<0.36\pi$, and in the vicinity of the RWL point $(\theta,T)=(\pi,T_{\rm RWL})$. The fine structure of the curve of $R=0$ in the vicinity of the RWL point is shown in Fig. 20. $R$ is negative elsewhere. The curve of $R=0$ near the $T$ axis is connected to the RWL region by the curve of $|RT^{3}|_{\rm min}$. There are two regions where $g^{\prime}=-g>0$ and $g_{11}^{\prime}=g_{11}>0$. There are also two regions where $g_{11}$ is negative. The curve determined by the condition $|n_{b0}|=B(T)=1/v_{\rm B}$ is the horizontal line which connects the upper zero point of $R$ at $\theta=0$ with the RWL point at $\theta=\pi$.
Fig. 19: The solid line represents the curve determined by the condition $R=0$ on the $\theta$-$T$ plane in the HRG model with EVE. The dash-dotted line represents the curve on which $|RT^{3}|$ has its minimum for the fixed value of $\theta$ in the region of $0.35<\theta<0.9914$ where the zero point of $R$ does not exist. The square is the RWL point $(\theta,T)=(\pi,T_{\rm RWL})$. The dotted (dash-dot-dotted) line represents the curve determined by the condition $g=0$ ($g_{11}=0$). The dashed line represents the curve determined by the condition $|n_{\rm b0}|=B(T)=1/v_{\rm B}$. Note that $|n_{\rm a0}|=|n_{\rm b0}|$ when $\mu$ is imaginary.
Fig. 20: The solid line represents the curve determined by the condition $R=0$ on the $\theta$-$T$ plane in the HRG model with EVE. The dash-dotted line represents the curve on which $|RT^{3}|$ has its minimum for the fixed value of $\theta$ in the region of $0.99\leq\theta<0.9914$ where the zero point of $R$ does not exists. The square is the RWL point $(\theta,T)=(\pi,T_{\rm RWL})$.
@@
Figure 22 shows the phase diagram determined by the condition $R=0$ on the $\mu^{2}$-$T$ plane when EVE is present. The curve of $R=0$ near the $T$ axis continues analytically from the real $\mu$ region to the imaginary one and is connected to the RWL region by the curve of $|RT^{3}|_{\rm min}$. The efficiency of the HRG model with/without EVE can be checked by calculating the scalar curvature $R$ by the LQCD calculation in the imaginary chemical potential region where there is no sign problem. In the imaginary $\mu$ region, $R$ diverges on the curves on which $g=0$. It seems that there is no counterpart of this singularity in the real $\mu$ region, since this divergence of $R$ due to the vanishing of $g$ is also present in the case without EVE.
@@
VI limiting temperature of baryon gas model
At $\theta=\pi$, $T_{\rm RWL}$ (=0.2103 GeV) can be regarded as the limiting temperature of the HRG model with EVE, or more precisely, the limiting temperature of the baryon gas model. This limitation is related with the sign change of the scalar curvature $R$. However, in QCD (not in the HRG model), it is expected that transition from the baryon gas to the quark matter occurs at $T=T_{\rm RW}$ which is slightly smaller than $T_{\rm RWL}$. On the other hand, in the region of real $\mu$, the critical point (CP) predicted by the LQCD calculation lies just below the right-lower curve of the $R=0$ criterion. This small disagreement indicates that the HRG model with EVE is not applicable beyond the critical point. As is the case of the RW transition, in the region of real $\mu$ in the QCD phase diagram, it is expected that there is true limiting temperature of the baryon gas, which is slightly lower than the temperature of the $R=0$ criterion in the HRG model with EVE. In this section, we investigate the limitation of the baryon gas model within the framework of the HRG model with EVE.
@@
Figure 23 shows $T$-dependence of $\chi_{2}^{\rm B}/T^{2}$ at $\mu=0$ in the HRG models with EVE and without EVE. In the HRG model without EVE, $\chi_{2}^{\rm B}/T^{2}$ increases monotonically as $T$ increases. In the HRG model with EVE, it has its maximum at $T_{\rm max}=0.195$ GeV. Due to the additional factor $T^{-3}$, $T_{\rm max}=0.195$ GeV is somewhat lower than $T_{\rm RWL}$. We see that the results in the HRG model with EVE are consistent with those in LQCD up to $T_{\rm max}$. This indicates that $T_{\rm max}$ is the limiting temperature of the baryon gas model.
@@
Figure 24 shows $T$-dependence of $\chi_{2}^{\rm B}/T^{2}$ at $\mu=0.6$ GeV in the HRG models with EVE and without EVE. Again, in the HRG model without EVE, $\chi_{2}^{\rm B}/T^{2}$ increases monotonically as $T$ increases. In the HRG model with EVE, it has two local maxima and one local minimum. The temperature of lower (upper) local maxima of $\chi_{2}^{\rm B}/T^{2}$ is somewhat lower than the temperature of lower (upper) local maximum of $\phi_{\gamma\gamma}/T_{\rm RWL}^{3}$. It is also seen that the lower local maximum of $\chi_{2}^{\rm B}/T^{2}$ is the global maximum. Hereafter, we regard the temperature of the global maximum of $\chi_{2}^{\rm B}/T^{2}$ as the limiting temperature of the baryon gas model.
In two dimensional conformal field theory, using the renormalization group method, it was shown that the effective degrees of freedom decrease as the energy scale of the system decreases Zamolodchikov (1986). It is called the c-theorem. Inversely, it may be natural that the effective degrees of freedom increase as the energy scale such as temperature increases. Without EVE, the effective degrees of freedom (EDOF) of the HRG model increases as $T$ increases. However, EVE reduces EDOF and the dimensionless quantity $\chi_{2}^{\rm B}/T^{2}$ has its maximum at $T_{\max}$ where the two effects are balanced. It is unnatural that EVE overcomes the effect of increase of temperature and makes $\chi_{2}^{\rm B}/T^{2}$ decrease as temperature increases. Hence, we regard the lower $T_{\rm max}$ as the limiting temperature of the HRG model with EVE.
@@
Figure 25 shows the $\mu$-dependence of $T_{\rm max}(\mu)$ ($T_{\rm min}(\mu)$) where $\chi_{2}^{\rm B}/T^{2}$ has its local maximum (minimum) for the fixed value of $\mu$ in the HRG model with EVE. We see that the critical point predicted by LQCD is located almost on the lower curve of $T_{\rm max}(\mu)$. This fact supports our assumption that the lower $T_{\rm max}(\mu)$ curve represents the limiting temperature of the baryon gas model. Above this temperature, quarks deconfine at least partially.
In Fig. 25 the $\mu$-dependence of $T_{\rm max}^{\prime}(\mu)$ ($T_{\rm min}^{\prime}(\mu)$) where $\phi_{\gamma\gamma}$ has its local maximum (minimum) for the fixed value of $\mu$ and the curve determined by the condition $n_{\rm b}=n_{\rm b0}/2=1/(2v_{\rm B})$ ($n_{\rm a}=n_{\rm a0}/2=1/(2v_{\rm B})$) are also shown. Note that $T_{\rm max}^{\prime}(\mu=0)=T_{\rm RWL}$. The lower (upper) curve of $T_{\rm max}^{\prime}$ almost coincides with the curve of $n_{\rm b}=1/(2v_{\rm B})$ ($n_{\rm a}=1/(2v_{\rm B})$) when $\mu$ is large. We see that the lower $T_{\rm max}(\mu)$ curve is just below the curve of $n_{\rm b}=1/(2v_{\rm B})$. This fact means that the limiting density $n_{\rm b}(\mu,T_{\rm max}(\mu))$ is smaller than $1/(2v_{\rm B})$. When $\mu$ is large, $n_{\rm B}\sim n_{\rm b}$. Hence, we obtain a simple empiric sufficient condition $n_{\rm B}>1/(2v_{\rm B})$ for quark deconfinement at large $\mu$, if we regard the lower $T_{\rm max}(\mu)$ as the limiting temperature of baryons.
Fig. 25: The dashed-dotted (solid) line shows the position of the local maximum of $\chi_{2}^{\rm B}/T^{2}$ ($\phi_{\gamma\gamma}/T_{\rm RWL}^{3}$) at fixed value of $\mu$ in the HRG model with EVE, while the dotted (dash-dot-dotted) line shows the position of its local minimum. The lower (upper) dashed line represents the curve determined by the condition $n_{b}=1/(2v_{\rm B}$) ($n_{a}=1/(2v_{\rm B})$). The meaning of the symbols is the same as in Fig. 14.
@@
Fig. 28: The phase diagram in $\mu$-$T$ plane. The square point is the cross point given by Eq. (70). The meaning of the lines and the other symbols is the same as in Figs. 14 and 25.
@@
It is very interesting that this cross point is very close to the critical point predicted by LQCD calculation. The crossover transition line is expected to end on the curve of the limitation of the baryon gas model. The endpoint is nothing but the critical point. Similarly, the curve of the $\chi_{2}^{\rm B}T/\varepsilon_{T}$ maximum in the HRG model should end at the cross point where the curve crosses the curve of the baryon gas limitation. Remember that, at $\mu=0.3$ GeV, the $\chi_{2}^{\rm B}T/\varepsilon_{T}$ maximum curve merges with the LQCD crossover transition line. Hence, this cross point may be regarded as an approximate critical point predicted by the HRG model with EVE, although the model has no explicit dynamics of the phase transition. Fig. 28 shows the phase diagram which is predicted by our HRG model with EVE, if we assume that the curve of the $\chi_{2}^{\rm B}/T^{2}$ global maximum gives the limiting temperature of baryons. Above the dashed-dotted line, quarks deconfine at least partially. Note also that, below the dashed line, $n_{\rm b}<1/(2v_{\rm B})$ in the HRG model with EVE.
VII Baryon radius dependence
In our model, we use the same baryon radius for all baryons. Hence, this baryon radius $r_{\rm B}$ is regarded as the averaged radius and its magnitude is uncertain. Here, we examine the $r_{\rm B}$-dependence of the results.
Fig. 29: The solid line shows the $r_{B}$-dependence of the Roberge-Weiss like temperature $T_{\rm RWL}$. The dashed line shows the result obtained by using the formula (71). The upper and lower dotted line show the maximum and minimum of the Roberge-Weiss temperature $T_{\rm RW}$ in the LQCD simulations Bonati et al. (2016); Cuteri et al. (2022); Bonati et al. (2019), respectively.
Fig. 29 shows the $r_{B}$-dependence of the Roberge-Weiss like temperature $T_{\rm RWL}$. It is seen that $T_{\rm RWL}$ decreases monotonically as $r_{\rm B}$ increases. The $r_{\rm B}$-dependence of $T_{\rm RWL}$ is well reproduced by the following formula.
$T_{\rm RWL}=0.1951{\rm GeV}/(r_{\rm B}/{\rm fm})^{1/3}$ (71)
Remember that $T_{\rm RWL}$ is determined by the condition $B(T_{\rm RWL})=1/v_{\rm B}$. It is easily seen that Eq. (71) is satisfied if $B(T)^{1/9}$ is proportional to $T$. In fact, in the region near $T\sim 0.2$ GeV, $B(T)^{1/9}/T$ is approximately constant. Hence, in this region, $B(T)$ follows the $T^{9}$-scaling law, the power of which is much greater than the one of the Stefan-Boltzmann limit, namely, 3. This strong power is induced by the succeeding generations of the hadron resonances. The higher the temperature becomes, the more significant the contributions from heavier hadrons become.
The Roberge-Weiss like temperature $T_{\rm RWL}$ is comparable to the so-called Lambda QCD $\Lambda_{\rm QCD}$. In the perturbative QCD (pQCD), $\Lambda_{\rm QCD}$ represents the lower energy limit for pQCD and the running coupling diverges at $\Lambda_{\rm QCD}$. On the contrary, $T_{\rm RWL}$represents the upper energy limit for the application of the HRG model with EVE and the thermodynamic quantities diverge at $T_{\rm RWL}$ when $\theta=(2k+1)\pi$ with any integer $k$.
Since the Roberge-Weiss like singularity in the HRG model with EVE is not observed in the LQCD simulation, transition from baryon matter to quark matter should take place before the temperature reaches $T_{\rm RWL}$. Hence, $T_{\rm RWL}$ should be larger than the Roberge Weiss temperature $T_{\rm RW}$ which lies in the region T=0.1940.213T=0.194-0.213 GeV  Bonati et al. (2016); Cuteri et al. (2022); Bonati et al. (2019). Hence, here, we regard $T=0.194$ GeV as the minimum value of the estimated $T_{\rm RW}$. Figure 29 shows that $r_{\rm B}\leq 1.02$ fm should be satisfied for the condition $T_{\rm RWL}\geq 0.194$ GeV.
On the other hand, since the RWL singularity is related to the baryon instability and the RW transition, it is not natural that $T_{\rm RWL}$ is much larger than $T_{\rm RW}$. Hence, here we consider the cases with $r_{\rm B}=0.7$ fm and 1.0 fm.
Fig. 30: The solid lines represent the curves determined by the condition $R=0$ on the $\mu$-$T$ plane in the HRG model with EVE when $r_{\rm B}=0.7$ fm. The dash-dotted line represents the curve on which $|RT^{3}|$ has its local minimum for the fixed value of $\mu$ in the region where 0.202 $\mu<0.544 GeV$ and $T<T_{\rm RWL}$. The lower (upper) dashed line represents the curve determined by the condition $n_{\rm b}=1/(2v_{\rm B})$ ($n_{\rm a}=1/(2v_{\rm B}$)). The meaning of the symbols is the same as in Fig. 14.
Fig. 31: The solid lines represent the curves determined by the condition $R=0$ on the $\mu$-$T$ plane in the HRG model with EVE when $r_{\rm B}=1.0$ fm. The dash-dotted line represents the curve on which $|RT^{3}|$ has its local minimum for the fixed value of $\mu$ in the region where 0.200 $\mu<0.495 GeV$ and $T<T_{\rm RWL}$. The lower (upper) dashed line represents the curve determined by the condition $n_{\rm b}=1/(2v_{\rm B})$ ($n_{\rm a}=1/(2v_{\rm B}$)). The meaning of the symbols is the same as in Fig. 14.
Figure 30 shows the $R=0$ condition in the HRG model with EVE when $r_{\rm B}=0.7$ fm. It is seen that the $R=0$ curve near the critical point is somewhat higher than the one in the model with $r_{\rm B}=0.8$ fm. Similarly, the lower $R=0$ curve in the small $\mu$ region is slightly higher than the LQCD crossover line. However, qualitative features of the diagram hardly changed from Fig. 15.
Figure 31 shows the $R=0$ condition when $r_{\rm B}=1.0$ fm. The $R=0$ curve near CP is somewhat lower than the one in the model with $r_{\rm B}=0.8$ fm and the CP locates almost on the curve. On the other hand, the lower $R=0$ curve in the small $\mu=0$ region is slightly lower than the crossover line. However, again, qualitative features of the diagram hardly changed from Fig. 15.
It should be noted that the $r_{\rm B}$-dependence of $T_{\rm RWL}$ in (71) is much weaker than $T_{\rm RWL}\sim 1/r_{\rm B}$ which is expected by the naive dimensional analyses. Thanks to the $T^{9}$-scaling of $B(T)$, $T_{\rm RWL}$ depends on $r_{\rm B}$ only weakly and the qualitative features of the results are robust under the moderate variations of $r_{\rm B}$.
$r_{\rm B}$  0.7 fm  0.8 fm  1.0 fm LQCD
$T_{\rm RWL}$ 0.220 0.210 0.195 0.1940.213
[…]
Table 1: Summary of the $r_{\rm B}$ dependence of the critical/limiting temperature shown in GeV. The $T_{3}$ is the temperature of the lower $R=0$ curve at $\mu=\bar{\mu}_{\rm CP}=0.6021$ GeV, the average value of the chemical potential of CP in the LQCD Shah et al. (2024). The LQCD reference value in the second row is the value of $T_{\rm RW}$ in Refs. Bonati et al. (2016); Cuteri et al. (2022); Bonati et al. (2019). The LQCD reference values in the third and fourth rows are the temperature of the crossover line at $\mu=0$ in Ref. Borsanyi et al. (2020), while the ones in the fifth and the sixth rows are the temperature $T_{\rm CP}$ of the critical point in Ref. Shah et al. (2024).
However, quantitative features of the results are somewhat changed under the variation of $r_{\rm B}$. Table 1 summarizes the critical or limiting temperature in the HRG model with EVE when $r_{\rm B}=0.7,0.8,1.0$ fm. In all cases, the condition $T_{\rm RWL}\geq 0.194$ GeV is satisfied. Note that $T_{\rm RWL}$ is equal to the temperature $T_{1}$ which is the temperature of the upper $R=0$ curve at $\mu=0$. We see that $T_{\rm CP}$ almost coincides with $T_{\rm max}(\bar{\mu}_{\rm CP})$ in the model with $r_{\rm B}=0.8$ fm, while it almost coincides with $T_{3}$ in the model with $r_{\rm B}=1.0$ fm. To eliminate the ambiguity of $r_{\rm B}$, more accurate determination of the Roberge-Weiss temperature $T_{\rm RW}$ is needed. However, since the critical point always locates below the curve of $n_{\rm b}=1/(2v_{\rm B})$ in all cases, the condition $n_{\rm B}>1/(2v_{\rm B})$ is an empiric sufficient condition for the deconfinement of quarks at large $\mu$.
Fig. 32: The solid lines represent the curves determined by the condition $R=0$ on the $\mu$-$T$ plane in the HRG model with EVE, when the baryon radius $r_{\rm b}=0.8$ fm and the antibaryon radius $r_{\rm a}=1.0$ fm. The dash-dotted line represents the curve on which $|RT^{3}|$ has its local minimum for the fixed value of $\mu$ in the region where 0.201 $\mu<0.525 GeV$ and $T<T_{\rm RWL}$. The lower (upper) dashed line represents the curve determined by the condition $n_{\rm b}=1/(2v_{\rm b})$ ($n_{\rm a}=1/(2v_{\rm a})$) where $v_{\rm b}={4\pi\over{3}}r_{\rm b}^{3}$ and $v_{\rm a}={4\pi\over{3}}r_{\rm a}^{3}$. The meaning of the symbols is the same as in Fig. 14.
It is expected that baryons are compressible to some extent before they collapse. Hence, at large $\mu$, the radius $r_{\rm b}$ of a baryon may be smaller than the radius $r_{\rm a}$ of an antibaryon. Figure 32 shows the $R=0$ condition in the HRG model with EVE when $r_{\rm b}=0.8$ fm and $r_{\rm a}=1.0$ fm. It is seen that the structure of the $R=0$ curve near the critical point is not hardly changed from Fig. 15, since the antibaryon contribution is negligible in that region. Of course, this setting is not appropriate at $\mu=0$, since it breaks the charge symmetry at that point. To preserve the charge symmetry at $\mu=0$, $r_{\rm B}$ should depend on $\mu$. However, in such a modification, the simple pressure forms (24) and (25) are not retained. Similarly, introduction of species-dependent $r_{\rm B}$ in the model is also very difficult in the present formulation. Hence, we postpone such modifications in the future research.
VIII Summary